# MA-RLHF: Reinforcement Learning from Human Feedback with Macro Actions

### v1

@@ 3.2.3 Connection to Previous Methods
MA-RLHF builds on and generalizes prior work in the RLHF literature by varying the length of macro actions. When the macro action length is set to $1$, MA-RLHF reduces to the standard token-level RLHF (Stiennon et al., 2020; Ouyang et al., 2022), operating as an MDP. Conversely, if we allow $|\omega_{\tau}|\rightarrow\infty$, then MA-RLHF converges toward methods like RLOO (Ahmadian et al., 2024) and REINFORCE (McGovern & Sutton, 1998), approximating a contextual bandit problem where decisions are made based on the entire sequence context. By varying the length of macro actions $|\omega_{\tau}|$, MA-RLHF provides a flexible framework that balances the granularity of action decisions. We provide further analysis on the impact of $|\omega_{\tau}|$ in §4.3.
@@ 4.1 Experimental Settings
Tasks and Datasets  We evaluate MA-RLHF on three different datasets for open-ended generation tasks: TL;DR (Stiennon et al., 2020) dataset for text summarization, Anthropic Helpful and Harmless (HH-RLHF) (Bai et al., 2022) for dialogue generation [^3] https://huggingface.co/datasets/Dahoas/full-hh-rlhf, and WebGPT Comparison (Nakano et al., 2021) for question answering. Additionally, we evaluate MA-RLHF on code generation using the APPS (Hendrycks et al., 2021) dataset. More details of the datasets can be found in Section B.1.
Base models and training details   For open-ended generation tasks, we use pre-trained Gemma-2B (Team et al., 2024) as our base model; we further adopt Gemma-7B and Gemma-2-27B to test the scaling trend. For the program synthesis task, we use CodeGemma-1.1-2B and CodeGemma-1.1-7B-it as our base models. The data split for SFT/RM/PPO and the hyperparameters used in SFT/RM/PPO stages are detailed in Section B.2. The implementation details of MA-PPO along with the code snippet can be found in Appendix E.
@@ 4.2 Main Results
Figure 3: Distribution of the RM scores for vanilla PPO and MA-PPO on TL;DR test set.
Figure 4: The win rates of MA-PPO against vanilla PPO, estimated by GPT-4 and Human. The left/middle/right figures are results for TL;DR, HH-RLHF and WebGPT Comparisons respectively.
@@
Moreover, Figure 3 highlights the distribution of RM scores, where MA-PPO consistently shifts towards higher RM sores compared to vanilla PPO. Further evaluation using GPT-4, given in the left figure of Figure 4, shows that MA-PPO achieves 78% and 86% win rate over vanilla PPO for the 2B and 7B models, respectively. Human evaluation gives similar result, where MA-PPO obtains win rates of 74% and 69%, and further confirming the effectiveness of macro actions. The final testing RM scores of MA-PPO and vanilla PPO are given in Table 2.
@@ 4.3.2 Ablation Study: Varying $n$ in MA-RLHF
GPT-4 Evaluation Analysis   As shown in Figure 7, setting $n=5$ strikes a good balance between relevance, coherence, consistency; it outperforms both smaller and larger values of $n$. These findings align with the Semi-MDP framework: increasing $n$ allows for better credit assignment and context retention, but excessive abstraction (e.g., $n=\infty$) sacrifices fine-grained control. Overall, moderate values of $n=5$ and $n=10$ provide the best trade-offs, highlighting the adaptability of MA-RLHF across tasks.
@@ 4.4 Generalization Probing in Macro Actions
Robustness on Rejection Sampling vs. Temperature   Best-of-$N$ (a.k.a, rejection sampling) (Touvron et al., 2023) enhances response quality by selecting the highest-reward response from $N$ samples generated by the policy model. We compare MA-PPO, SFT, and vanilla PPO using the best-of-$N$ sampling across various temperatures $T\in\{0.2,0.4,0.6,0.8,1.0,1.2\}$ and sample sizes $N\in\{4,8,16,32\}$. As shown in Figure 8, best-of-$N$ sampling improves RM scores for all methods, with performance increasing as $N$ grows. We observe that SFT and vanilla PPO are sensitive to temperature variations, requiring specific adjustments to achieve optimal results. In contrast, MA-PPO demonstrates robustness in sampling temperature, it consistently delivers the best performance at $T=1.2$ and shows consistent improvement across all tested temperatures. Moreover, MA-PPO maintains stable performance across varying temperature settings, as further detailed in Section D.3, highlighting its robustness and generalization capabilities under different sampling temperatures.
Figure 9: Evaluation results for vanilla PPO and MA-PPO on Gemma-2-27B using the TL;DR dataset. Left: RM scores on validation set. Mid: Distribution of RM scores for vanilla PPO and MA-PPO. Right: Scaling trending on TL;DR dataset across 2B, 7B, and 27B model size, showing RM scores, GPT-4 evaluation, human evaluation results.
@@ 4.5 Additional Analysis
Impact on $L_{2}$-norm of Advantage and Q Values   We present the $L_{2}$-norm of both the advantage and Q-values for MA-PPO and vanilla PPO during training in Figure 12. The advantage function, which reflects the difference between the expected return (Q-value) and the baseline, plays a critical role in guiding policy optimization. A lower $L_{2}$-norm of both the advantage and Q-values suggests more stable and less noisy policy updates, likely contributing to faster convergence observed in §4.2.
@@
Extended Experiments: Program Synthesis   We further assess the effectiveness of MA-PPO on the program synthesis task. Following (Shojaee et al., 2023; Liu et al., 2023), we utilize the compiler signal as the final reward; see Section B.5 for implementation details. We compare the performance of MA-PPO and vanilla PPO using the pass@k (k=1, 5) metric Chen et al. (2021) on the 5k test set from APPS dataset (Hendrycks et al., 2021). As shown in Figure 12, MA-PPO significantly outperforms vanilla PPO across both pass@1 and pass@5 metrics, with more pronounced improvements as model size scales. Notably, for the 7B model, MA-PPO achieves an improvement of +35% in pass@1 and +30% in pass@5 over vanilla PPO, demonstrating the effectiveness of our approach in code generation tasks.
@@ 5 Related Work
LLM Alignment   RLHF have shown impressive success in aligning LLMs with human preferences through multi-stage training, including SFT, RM, and RL fine-tuning (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). Recent research has explored optimization methods for RL in LLMs, employing both online (Ahmadian et al., 2024; Farebrother et al., 2024; Shen et al., 2024; Chakraborty et al., 2024) and offline RL algorithms (Snell et al., 2023; Hu et al., 2023; Yu et al., 2024) to address training instability and improve efficiency (Tang et al., 2024). Improvements to RM learning have been proposed, such as parameter scaling (Gao et al., 2023), fine-grained reward (Wu et al., 2023), tool use (Li et al., 2024), and model merging (Ramé et al., 2024; Rame et al., 2024). Alternatively, direct policy optimization (Rafailov et al., 2024; Ethayarajh et al., 2024; Gheshlaghi Azar et al., 2023; Rosset et al., 2024) has emerged as a promising approach, bypassing the instability of RL while directly aligning models to human preferences. In this paper, we enhance the RLHF action space by integrating macro actions, a well-established concept in RL(Sutton et al., 1999; Mann & Mannor, 2014).
Macro Action in RL   Macro actions introduce temporal abstraction in RL by grouping sequences of primitive actions, reducing decision complexity and improving long-horizon credit assignment (Precup et al., 1997; Hauskrecht et al., 2013; Sutton et al., 1999; Pignatelli et al., 2024; Machado et al., 2023). This method has demonstrated its utility in speeding up convergence and stabilizing policy updates in various domains (Mann & Mannor, 2014; Solway et al., 2014). Our work applies macro actions to RLHF in LLM training, leveraging this structure to enhance scalability and optimize credit assignment over extended sequences.
@@ 6 Conclusion and Future Work
In this paper, we introduced MA-RLHF, a novel framework that incorporates macro actions into RLHF to enhance the alignment of LLMs with human preferences. Our approach demonstrates consistent improvements across multiple tasks, including summarization, dialogue generation, question answering, and program synthesis. Notably, MA-RLHF achieves 1.7x to 2x faster convergence in reward scores without incurring additional computational overhead, showing robust scalability across model sizes ranging from 2B to 27B parameters. It is promising to explore more fine-grained or learnable macro action strategies for future research.
@@ Appendix A Limitations
While our work demonstrates the effectiveness of MA-RLHF across multiple tasks, there are several limitations that leave room for future improvements. First, our macro action termination methods are rule-based, including linguistics- or perplexity-driven approaches; future research could explore more complex or learnable termination strategies to further enhance performance. Second, our experiments are conducted using models with up to 27B parameters; exploring more advanced models, such as LLaMA 3.1 405B (Dubey et al., 2024) or other state-of-the-art architectures, may provide additional insights into the scalability of MA-RLHF. Lastly, although we observe significant improvements in training efficiency and convergence speed, further investigation into the trade-offs between training stability and performance under diverse real-world conditions is necessary. Addressing these limitations will pave the way for more robust applications of MA-RLHF.
@@ B.5 Training Settings of Program Synthesis
$R(x)=\left\{\begin{aligned} &-0.3+1.3\cdot\frac{N_{\text{pass}}}{N_{\text{pass}}+N_{\text{fail}}},&\text{if }x\text{ successfully compiled.}\\ &-0.6,&\text{if }x\text{ received runtime error.}\\ &-1.0,&\text{if }x\text{ received compile rrror.}\end{aligned}\right.$
where $x$ represents the code snippet generated by the policy model.
@@ C.1 Results of Dialogue Generation
Figure 15: Distribution of test RM scores for vanilla PPO and MA-PPO on Gemma-2B using the HH-RLHF dataset.
@@ C.2 Results of Question Answering
Figure 17: Distribution of test RM scores for vanilla PPO and MA-PPO on WebGPT dataset.
@@ D.1 Value Function Estimation of Macro Action
D.2 Impact of RLHF on Reward Score Distribution
Figure 20: RM score shifting pattern after RLHF training. Left presents the RM score of best of 8 sampling on vanilla PPO compared to the vanilla PPO. Right presents the RM score of best of 8 sampling on MA-PPO compared to the MA-PPO.

### v2

@@ 3.2.3 Connection to Previous Methods
MA-RLHF builds on and generalizes prior work in the RLHF literature by varying the length of macro actions. When the macro action length is set to $1$, MA-RLHF reduces to the standard token-level RLHF (Stiennon et al., 2020; Ouyang et al., 2022), operating as an MDP. Conversely, if we allow $|\omega_{\tau}|\rightarrow\infty$, then MA-RLHF converges toward methods like RLOO (Ahmadian et al., 2024), REINFORCE (Williams, 1992; Sutton et al., 1999a), and GRPO (Shao et al., 2024), approximating a contextual bandit problem where decisions are made based on the entire sequence context. By varying the length of macro actions $|\omega_{\tau}|$, MA-RLHF provides a flexible framework that balances the granularity of action decisions. We provide further analysis on the impact of $|\omega_{\tau}|$ in §4.3.
@@ 4.1 Experimental Settings
Tasks and Datasets  We evaluate MA-RLHF on three different datasets for open-ended generation tasks: TL;DR (Stiennon et al., 2020) dataset for text summarization, Anthropic Helpful and Harmless (HH-RLHF) (Bai et al., 2022) for dialogue generation [^3] https://huggingface.co/datasets/Dahoas/full-hh-rlhf, and WebGPT Comparison (Nakano et al., 2021) for question answering. Additionally, we evaluate MA-RLHF on code generation using the APPS (Hendrycks et al., 2021) dataset. More details can be found in Section B.1.
Base Models and Training Details   For open-ended generation tasks, we use pre-trained Gemma-2B (Team et al., 2024) as our base model; we further adopt Gemma-7B and Gemma-2-27B to test the scaling trend. For the program synthesis task, we use CodeGemma-1.1-2B and CodeGemma-1.1-7B-it as our base models. The data split for SFT / RM / PPO and the hyperparameters used in SFT / RM / PPO stages are detailed in Section B.2. The implementation details of MA-PPO can be found in Appendix E.
@@ 4.2 Main Results
Figure 3: RM score distribution for PPO and MA-PPO (2B) at final steps (4.6k) on TL;DR.
Figure 4: Win rates of MA-PPO against vanilla PPO on TL;DR (left), HH-RLHF (middle) and WebGPT Comparisons (right), estimated by GPT-4 and Human.
@@
Moreover, Figure 3 highlights the distribution of RM scores, where MA-PPO consistently shifts towards higher RM sores compared to vanilla PPO. Further evaluation using GPT-4, given in the left figure of Figure 4, shows that MA-PPO achieves 78% and 86% win rate over vanilla PPO for the 2B and 7B models, respectively. Human evaluation gives similar results, where MA-PPO obtains win rates of 74% and 69%, further demonstrating the effectiveness of macro actions. The final testing RM scores of MA-PPO and vanilla PPO are given in Table 2.
@@ 4.3.2 Ablation Study: Varying $n$ in MA-RLHF
GPT-4 Evaluation Analysis   As shown in Figure 7, setting $n=5$ strikes a good balance between relevance, coherence, consistency; it outperforms both smaller and larger values of $n$. These findings align with the semi-MDP framework: increasing $n$ allows for better credit assignment and context retention, but excessive abstraction (e.g., $n=\infty$) sacrifices fine-grained control. Overall, moderate values of $n=5$ and $n=10$ provide the best trade-offs, highlighting the adaptability across tasks.
@@ 4.4 Generalization Probing in Macro Actions
Robustness on Rejection Sampling vs. Temperature   Best-of-$N$ (a.k.a, rejection sampling) (Touvron et al., 2023) enhances response quality by selecting the highest-reward response from $N$ samples generated by the policy model. We compare MA-PPO, SFT, and vanilla PPO using the best-of-$N$ sampling across various temperatures $T\in\{0.2,0.4,0.6,0.8,1.0,1.2\}$ and sample sizes $N\in\{4,8,16,32\}$. As shown in Figure 8, best-of-$N$ sampling improves RM scores for all methods, with performance increasing as $N$ grows. We observe that SFT and vanilla PPO are sensitive to temperature variations, requiring specific adjustments to achieve optimal results. In contrast, MA-PPO demonstrates robustness in sampling temperature, it consistently delivers the best performance at $T=1.2$ and shows consistent improvement across all tested temperatures. Moreover, MA-PPO maintains stable performance across varying temperature settings, as detailed in Section D.4, highlighting its robustness and generalization capabilities under different sampling temperatures.
Figure 9: Evaluation results for vanilla PPO and MA-PPO on Gemma-2-27B using the TL;DR dataset. Left: RM scores on validation set. Mid: Distribution of RM scores for vanilla PPO and MA-PPO (27B) at final steps (4.6k). Right: Scaling trending on TL;DR dataset across 2B, 7B, and 27B model size, showing RM scores, GPT-4 evaluation, human evaluation results.
@@ 4.5 Additional Analysis
Impact on $L_{2}$-Norm of Advantage and Q Values   We present the $L_{2}$-norm of both the advantage and Q-values for MA-PPO and vanilla PPO during training in Figure 12. The advantage function, which reflects the difference between the expected return (Q-value) and the baseline, is critical in guiding policy optimization. A lower $L_{2}$-norm of both the advantage and Q-values suggests more stable and less noisy policy updates, likely contributing to faster learning speed observed in §4.2.
@@
Extended Experiments: Code Generation   We further assess the effectiveness of MA-PPO on the code generation task. Following Shojaee et al. (2023); Liu et al. (2023), we utilize the compiler signal as the final reward; see Section B.5 for implementation details. We compare the performance of MA-PPO and vanilla PPO using the pass@k (k=1, 5) metric (Chen et al., 2021) on the 5k test set of the APPS dataset (Hendrycks et al., 2021). As shown in Figure 12, MA-PPO significantly outperforms vanilla PPO in both pass @ 1 and pass @ 5 metrics, with more pronounced improvements as model size scales. Notably, for the 7B model, MA-PPO achieves an improvement of +35% in pass@1 and +30% in pass@5 over vanilla PPO, demonstrating the effectiveness of our approach in code generation tasks.
@@ 5 Related Work
LLM Alignment   RLHF have shown impressive success in aligning LLMs with human preferences through multi-stage training, including SFT, RM, and RL fine-tuning (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Sun et al., 2025). Recent research has explored optimization methods for RL in LLMs, employing both online (Ahmadian et al., 2024; Farebrother et al., 2024; Shen et al., 2024; Chakraborty et al., 2024; Shao et al., 2024) and offline RL algorithms (Snell et al., 2023; Hu et al., 2023; Yu et al., 2024) to address training instability, improve efficiency (Tang et al., 2024) and diversity (Sun et al., 2025). Improvements to RM learning have been proposed, such as parameter scaling (Gao et al., 2023), fine-grained reward (Wu et al., 2023), tool use (Li et al., 2024), and model merging (Ramé et al., 2024; Rame et al., 2024). Alternatively, direct policy optimization (Rafailov et al., 2024; Ethayarajh et al., 2024; Gheshlaghi Azar et al., 2023; Rosset et al., 2024) has emerged as a promising approach, bypassing the instability of RL while directly aligning models to human preferences. In this paper, we enhance the RLHF action space by integrating macro actions, a well-established concept in RL (Sutton et al., 1999b; Mann & Mannor, 2014).
Macro Action in RL   Macro actions introduce temporal abstraction in RL by grouping sequences of primitive actions, reducing decision complexity and improving long-horizon credit assignment (Precup et al., 1997; Hauskrecht et al., 2013; Sutton et al., 1999b; Pignatelli et al., 2024; Machado et al., 2023a). This method has demonstrated its utility in speeding up convergence and stabilizing policy updates in various domains (Mann & Mannor, 2014; Solway et al., 2014). Our work applies macro actions to RLHF in LLM training, leveraging this structure to enhance scalability and optimize credit assignment over extended sequences.
@@ 6 Conclusion and Future Work
In this paper, we introduced MA-RLHF, a novel framework that incorporates macro actions into RLHF to enhance the alignment of LLMs with human preferences. Our approach demonstrates consistent improvements across multiple tasks, including summarization, dialogue generation, question answering, and code generation. Notably, MA-RLHF achieves parity with vanilla RLHF 1.7x to 2x faster in reward scores without incurring additional computational overhead, showing robust scalability across model sizes ranging from 2B to 27B parameters. It is promising to explore MA-RLHF in complex step-by-step reasoning tasks for future research.
Reproducibility Statement
We are committed to ensuring the reproducibility of the experiments presented in Section 4. To this end, we make the source code and model checkpoints publicly available at https://github.com/ernie-research/MA-RLHF. The detailed source code for training and evaluating both the conventional RLHF and our proposed MA-RLHF approach is included in the supplementary materials. We believe that these efforts will enable researchers to rigorously verify our findings and build upon our work.
Acknowledgments
We would like to express our gratitude to the anonymous reviewers for their insightful and constructive feedback.
@@ Appendix A Limitations
While our work demonstrates the effectiveness of MA-RLHF across multiple tasks, there are several limitations that leave room for future improvements. In our implementation, we apply the identical action / vocabulary space as pretrained LLMs, considering the fact that defining macro actions as one options (e.g., one macro action per $n$-gram) would require re-architecting the LLM’s vocabulary and retraining the model, which is computationally infeasible. Meanwhile, our macro action termination methods are rule-based, including linguistics- or perplexity-driven approaches; future research could explore more complex or learnable termination strategies to further enhance performance. Furthermore, regarding the generalization of MA-RLHF, our experiments are conducted using models with up to 27B parameters; exploring more advanced models, such as LLaMA 3.1 405B (Dubey et al., 2024) or other state-of-the-art architectures and tasks (e.g., mathematical and complex reasoning), may provide additional insights into the scalability of MA-RLHF. Lastly, although we observe significant improvements in training efficiency, further investigation into the trade-offs between training stability and performance under diverse real-world conditions is necessary. Addressing these limitations will pave the way for more robust applications of MA-RLHF.
@@ B.5 Training Settings of Program Synthesis
$R(x,y)=\left\{\begin{aligned} &-0.3+1.3\cdot\frac{N_{\text{pass}}}{N_{\text{pass}}+N_{\text{fail}}},&\text{if }y\text{ successfully compiled.}\\ &-0.6,&\text{if }y\text{ received runtime error.}\\ &-1.0,&\text{if }y\text{ received compile rrror.}\end{aligned}\right.$
where $x$ represents the prompt, and $y$ represents the code snippet generated by the policy model.
@@ C.1 Results of Dialogue Generation
Figure 15: Distribution of test RM scores for vanilla PPO and MA-PPO (2B) at final steps (5.6k) on the HH-RLHF dataset.
@@ C.2 Results of Question Answering
Figure 17: Distribution of test RM scores for vanilla PPO and MA-PPO (2B) at final steps (3.2k) on WebGPT dataset.
@@
Figure 18: Test RM scores of SFT model, vanilla PPO, MA-PPO, and baselines: DPO and RLOO on TL;DR and HH-RLHF datasets.
Method
RM Score
(TL;DR)
RM Score
(HH-RLHF)
SFT -0.64 0.13
DPO 0.03 0.64
RLOO 0.81 -
PPO 0.83 1.31
MA-PPO (n=5) 1.40 1.55
Figure 19: Win rates of DPO and RLOO against PPO and MA-PPO on TL;DR and HH-RLHF estimated by GPT-4.
C.3 Comparing with Additional Baselines
In this section, we compare MA-PPO with two additional baselines: DPO (Rafailov et al., 2024) and RLOO (Ahmadian et al., 2024) on Gemma-2B model. Both of the methods are implemented with Deepspeed-Chat. Specifically, DPO models are trained on TL;DR and HH-RLHF datasets, with the same data split as we used when training PPO. RLOO model is trained on TL;DR dataset only, with the same policy and reward model initialization as PPO. For the training details of DPO, the learning rate is set to 2e-7, with $\beta=0.1$ for TL;DR and $\beta=0.01$ for HH-RLHF. The policy and reference models are initialized using the same SFT model as in PPO. For RLOO, the learning rate for the policy model is set to 1.5e-5, and the number of online samples is $K=4$. All other hyperparameters are kept consistent with PPO.
We demonstrate the results evaluated by reward model score in Figure 19, and win rates estimated by GPT-4 in Figure 19. On TL;DR dataset, DPO fails to gain improvement compared to PPO and MA-PPO, while RLOO achieves similar performance compared to PPO, but outperformed by MA-PPO. On HH-RLHF dataset, DPO exhibits superior performance than PPO but still underperforms the MA-PPO.
Table 6: Test RM scores of Llama-3.2-3B models on TL;DR dataset.
Method RM Score (TL;DR)
SFT 2.38
PPO 3.33
MA-PPO (n=5) 3.96
C.4 Experiments on Llama-3.2-3B
We conduct experiments on Llama-3.2-3B model to validate the generalizability of our method across different model families. The experiments are conducted on TL;DR dataset, following the same data split as Gemma-2B. We set the learning rates of actor and critic to 5e-6 and 1e-5, and the KL coefficient is set to 0.1. Table 6 demonstrate the results evaluated by RM score, we show MA-PPO still remarkably outperforms vanilla PPO. Using GPT-4 to assess the win rate, MA-PPO obtains 61% win, 4% tie and 34% loss rate compared against PPO. These results prove the generalizability of our method.
@@ D.1 Value Function Estimation of Macro Action
Table 7: Pass@$1$ metric evaluated when applying different termination conditions on APPS dataset.
Dataset Termination RM Score
GPT-4 Win Rate
(v.s. PPO)
TL;DR Fixed 5-gram 1.40 78%
Parsing 1.37 78%
PPL 1.27 72%
HH-RLHF Fixed 5-gram 1.55 58%
Parsing 1.64 62%
Table 8: Test RM scores and GPT-4 win rates when applying different termination conditions on TL;DR and HH-RLHF datasets.
Termination Fixed 10-gram Parsing PPL
pass@1 Inter. 3.25 3.17 3.04
Intro. 16.56 17.05 16.36
Comp. 0.94 1.24 0.80
All 5.45 5.56 5.26
D.2 Termination Conditions on Different Tasks
In this section, we analysis the effectiveness of termination conditions on TL;DR, HH-RLHF, and APPS datasets. When implementing parsing-based termination condition on APPS dataset, we use a programming-language-based parser. [^4] RedBaronhttps://github.com/PyCQA/redbaron The results of TL;DR and HH-RLHF datasets are shown in Table 8 and Table 8. We can notice that parsing-based termination condition performs well on the HH-RLHF tasks, with higher RM score and win rate than fixed 5-gram based termination condition. While on the TL;DR dataset, parsing-based termination condition also achieves excellent performance compared to fixed 5-gram termination condition. On APPS dataset, parsing-based termination condition achieves the best results, except for the interview level task. These results demonstrate that construct macro action with linguistic information indeed brings performance gain to MA-PPO.
D.3 Impact of RLHF on Reward Score Distribution
Figure 22: RM score shifting pattern after RLHF training. Left presents the RM score of best of 8 sampling on vanilla PPO compared to the vanilla PPO. Mid Left presents the RM score of best of 8 sampling on MA-PPO compared to the MA-PPO. Mid Right presents the RM score of MA-PPO ($n=5$) compared to the vanilla PPO model. Right presents the RM scores of MA-PPO ($n=\infty$) compared to the vanilla PPO model.